What is the area in yd2 If the length is 10yd and the width is 36ft
Answer by jim_thompson5910(35256)

3 ft = 1 yd

12*3 ft = 12*1 yd

36 ft = 12 yd

-------------------------------------------------------

A = L*W

A = 10*12

A = 120

-------------------------------------------------------

The area is 120 square yards
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------